OpenStudy (tanabugg):

The pair of points (6,y) and (10, -1) lies on a line with a slope of 1/4. What's the value of y? a. -5 b. -2 c. 2 d. 5

OpenStudy (harman.singh):

Use the set of coordinates (10,-1) & the value for gradient 1/4 to find equation of the line in the form of y=mx+c . Once you have the equation, sub in the x-value of 6 and solve for y. That will give you the answer to your question

OpenStudy (tanabugg):

i know I probably sound unappreciative but do you know the answer?? @harman.singh

OpenStudy (harman.singh):

I do know how to find the answer but I think its better if you do it. You can ask me as many times you like if you get stuck. Here is more information: \[y-y _{1}=m(x-x_{1})\] \[y+1=m(x-10)\] Can you try finding the equation from here?

OpenStudy (tanabugg):

What is m stand for?

OpenStudy (harman.singh):

m stands for the gradient or the slope. It is given to you in the question (1/4)

OpenStudy (harman.singh):

So you can sub in m=1/4 into the equation y+1=m(x−10) y+1=1/4(x-10) and continue from there..